# leftmargin=0pt, rightmargin=0pt, vskip=0pt

### v1

@@ 5.3. Generalization to Multiple Sensors
The embedding interpolation methodology naturally extends to higher dimensions for scenarios involving more sensors. In such cases, the sensor space is divided into non-overlapping simplices (e.g., tetrahedra in 3D) using the generalized form of Delaunay triangulation, known as Delaunay tessellation (Watson, 1981), based on the anchor points. Figure 7 illustrates that a space containing three sensors and eight anchors can be subdivided into five tetrahedra. The embedding $E_{p}$ for a new sensor reading $P$ can then be computed as a linear combination of the embeddings of the anchors $A_{1}$, $A_{2}$, $A_{4}$, and $A_{5}$. The generalized process for calculating embeddings is outlined below:
@@ 6   Sensor-Aware Visual Production
Existing image generation models, such as Stable Diffusion, demonstrate the capability to generate visual content from CLIP embeddings (Face, 2023), and these images can be transformed into 3D scenes using text-to-3D and image-to-3D models (Tang et al., 2023; Poole et al., 2022; Liu et al., 2023). However, sensor data is highly context-dependent and features volumetric and dynamic characteristics that current embedding-to-image and 3D generation models are not equipped to handle.
To overcome these challenges, we leverage the foundation models to incorporate domain knowledge and sensor schema to enrich sensor data to produce 2D visual representations. We further employ a customized 3DGS generation model to encode spatial density, enabling the visualization of sensor data in 3D space. Additionally, to support low-latency visualization of dynamic sensor data, we design latent reuse caches at each step, enabling just-in-time AR presentations (Figure 8).
@@ 6.2. AR Production with Foundation Models
From the CLIP embeddings and sensor context, we produce 2D visualizations. Building on this foundation, we extend these images into 3D AR content. Leveraging 3DGS (Kerbl et al., 2023; Liu et al., 2023; Tang et al., 2023), we generate spatially-aware visualizations that capture the volumetric and spatial aspects of sensor data, such as temperature distributions within a space or air quality surrounding a street.
Traditional 3D content generation often relies on mesh-based models that primarily represent surface structures while leaving interiors void. While effective for certain applications, this approach is inadequate for sensor data, which often encapsulates volumetric phenomena. For example, environmental readings like temperature or humidity demand a representation that is volumetrically consistent rather than a hollow shell. To overcome this limitation, we propose a novel approach to 3D generation that incorporates the sensor embeddings and manifestations defined earlier, producing volumetric 3D models that accurately reflect the spatial context of sensor data.
First, our method integrates sensor embeddings and manifestations into the 3D generation pipeline by adapting the DreamGaussian (Tang et al., 2023) framework. Instead of standard inputs and outputs, we use sensor-specific embeddings and 2D images to generate a volumetric representation with 3D Gaussian Splatting. This preserves the inherent spatial characteristics of the sensor data, enabling more faithful visualizations.
To simulate volumetric features of sensor measurement, we introduce a novel volumetric generation that encodes density information into the generation. Specifically, we apply random dropout to Gaussian points to create transparent areas. This dropout enables the generation model to fill both internal and external spaces with Gaussian points based on multi-angle observations. The dropout rate is determined by the sensor schema and readings through a density encoder, ensuring alignment with the sensor data’s characteristics. This design ensures the generated AR content’s interior is fully populated with Gaussian points.
@@
Embedding diverse sensor data into visual spaces presents significant challenges, particularly in extreme conditions. Our proposed FM-assisted pipeline addresses these issues by leveraging large language models (LLMs) and schema generation. This approach tackles the inherent variability of sensor data, which stems from different sources and contexts, resulting in unique modalities. The pipeline effectively manages challenges such as diverse sensor types, context-dependent readings, and sparse data, offering an innovative solution for processing and embedding sensor data into visual or semantic spaces.
LLMs offer a robust solution for addressing the complexity and diversity of sensor data. These models excel in processing and understanding the intricate relationships between various data modalities. A key feature of our pipeline is the creation of a sensor schema that includes anchors specific to each sensor type and context. This schema is constructed using a combination of data collection, domain knowledge, and the capabilities of LLMs, enabling zero-shot interpretation by generating meaningful visual manifestations without prior task-specific training.
Our pipeline is specifically designed to handle sensor data using zero-shot learning capabilities, which allows for the automatic processing of new sensor inputs. We construct a schema for each sensor type that anchors the sensor data to its contextual meaning, leveraging LLMs for schema refinement. Through the use of prompt engineering, LLMs automatically generate sensor metadata descriptions, which are then visualized. Additionally, advanced image generation models are employed to create detailed and accurate visual constraints, ensuring consistency with the defined sensor schema and manifestation.
This approach enables the dynamic production of visual content for new sensor inputs without manual intervention, offering a scalable solution for handling the diverse modalities of sensor data. Through the integration of LLMs and innovative schema generation techniques, our pipeline effectively bridges the gap between raw sensor data and meaningful visual representations, facilitating easier interpretation and utilization of sensor data across various applications.
@@ 6.3. Latent Reuse for Fast Generation
The generation of both images and 3D Gaussian Splatting is computationally expensive. Even with high-end GPUs like the Nvidia RTX 4090, creating a single image can take over 5 seconds per frame, while generating a 3D scene with DreamGaussian (Tang et al., 2023) may require up to 50 seconds. These time demands significantly impact the practicality of sensor visualizations, especially when dealing with diverse and rapidly changing data. To mitigate this, we designed a latent reuse caching mechanism that accelerates the generation of both images and 3D scenes, achieving a 7$\times$ reduction in processing time without compromising output quality.
@@ 7.3. Evaluations of Visualization Quality
To assess the quality of generated content, we conducted a survey with 565 participants recruited through social media and forums, a total of 485 valid responses of which 91.2% had prior experience with sensor technologies, enhancing their evaluation accuracy. The survey aimed to compare Vivar with state-of-the-art image generation models based on three criteria, each rated on a scale from 1 to 5, where 1 is bad, 3 is okay, and 5 is awesome:
@@
Figure 10. Examples from our solution and compared models in two contexts: human emotion representation and temperature in a shared environment.
@@
    Ambient Environment: Participants imagined being in a room with temperature sensors reflecting temperature variations (measured in °C). The visualizations were intended to reflect these changes.
@@
    Emotion: Participants viewed images depicting a person experiencing a range of emotions. The goal was for the models to effectively capture and differentiate these emotional states.
@@
    Art Entertainment: Participants imagined listening to different types of music, from soft to noisy, expecting the visualizations to correspond to these auditory variations.
Figure 11. Left: Performance comparison of models across different contexts. Right: user considerations while doing the survey.
Results and Analysis
Table 2 summarizes the average scores and standard deviations for each model, with the highest scores for each metric highlighted in green and the lowest in red. Our proposed system, Vivar, achieved the highest scores in coherence (3.33) and faithfulness (3.20), indicating that participants rated it above okay and closer to awesome in these categories. This superior coherence is attributed to the continuity provided by embedding blending, resulting in logically consistent image generation. The high faithfulness score reflects the effectiveness of our manifestation generation in accurately representing sensor data. Although Vivar ranked second in sensitivity (3.22), it was close to DALL-E 3’s score (3.27).
@@
ATitan 3.11 $\pm$ 0.97 3.18 $\pm$ 1.01 3.11 $\pm$ 1.16 3.13
DallE3 3.10 $\pm$ 1.00 3.16 $\pm$ 1.02 3.27 $\pm$ 1.05 3.18
SDE 3.03 $\pm$ 1.07 3.07 $\pm$ 1.08 3.04 $\pm$ 1.24 3.05
SDP 3.03 $\pm$ 1.03 2.93 $\pm$ 1.03 2.93 $\pm$ 1.07 2.96
Vivar (Ours) 3.33 $\pm$ 0.97 3.20 $\pm$ 0.86 3.22 $\pm$ 1.05 3.25
Table 2. Mean and Standard Deviation of Coherence, Faithfulness, and Sensitivity for Each Model, along with the Overall Score calculated using the weights from user considerations. The models include ATitan: Amazon Titan, DallE3: DALL-E 3, SDE: Stable Diffusion with Embedding Input, SDP: Stable Diffusion with Prompt Input, and Vivar.
Ablation Study: SDE and SDP are subcomponents of Vivar. SDP uses conventional prompt-to-image generation, while SDE applies embedding blending without the manifestation component. As shown in Table 2, SDP performed worst across all metrics. Introducing embedding blending (SDE) significantly improved faithfulness and sensitivity, demonstrating that embedding blending enhances the system’s ability to generate more accurate and faithful visual representations compared to prompt-based approaches.
@@
The results are summarized in Table 2, allowing direct comparison of models based on user-prioritized criteria. Vivar achieves the highest performance. Notably, Vivar attains an overall score of 3.25, outperforming the state-of-the-art models DallE3 and ATitan by approximately 2.2% and 3.8%, respectively. This improvement demonstrates the effectiveness of our approach in aligning with user preferences.
@@ 9   Conclusion
In this work, we addressed the challenge of visualizing multi-modal and context-dependent sensor data by introducing a novel cross-modal embedding method that blends sensor readings into a pre-trained visual embedding space, guided by the shading principle. By incorporating diverse manifestations and 3DGS, our approach enables accurate, consistent visual representations with smooth transitions between data points, maintaining semantic coherence. The inclusion of a latent reuse cache further enhances adaptability to dynamic sensor readings. Through evaluations involving over 450 participants, we demonstrated that our solution outperforms state-of-the-art visual generation models in consistency and representativeness, making it better suited to sensor data visualization. A user study with domain experts further validated the system’s accuracy, reliability, and practical applicability across diverse downstream tasks.
This work establishes a foundation for intuitive, coherent, and unbiased sensor data visualization by bridging domains. Drawing inspiration from graphics and shading principles, we combine large pre-trained models with explainable transformations, opening new avenues for advancing cross-modal embedding techniques.

### v2

@@ 5.3. Generalization to Multiple Sensors
Our approach can be naturally extended to higher dimensions for scenarios involving more sensors. In such cases, the sensor space is divided into non-overlapping simplices (e.g., tetrahedra in 3D) using the generalized form of Delaunay triangulation, known as Delaunay tessellation (Watson, 1981), based on the anchor points. Figure 8 illustrates that a space containing three sensors and eight anchors can be subdivided into five tetrahedra. The embedding $E_{p}$ for a new sensor reading $P$ can then be computed as a linear combination of the embeddings of the anchors $A_{1}$, $A_{2}$, $A_{4}$, and $A_{5}$. The generalized process for calculating embeddings is outlined below:
@@ 6   Sensor-Aware Visual Production
Existing image generation models, such as Stable Diffusion, demonstrate the capability to generate visual content from CLIP embeddings (Face, 2023), and these images can be further used to generate into 3D scenes using text-to-3D and image-to-3D models (Tang et al., 2023; Poole et al., 2022; Liu et al., 2023). However, sensor data is highly context-dependent and features volumetric and dynamic characteristics that current embedding-to-image and 3D generation models are not equipped to handle.
To overcome these challenges, we leverage the foundation models to incorporate domain knowledge and sensor schema to enrich sensor data to produce 2D visual representations. We further employ a customized 3D Gaussian Splatting (3DGS) model to encode spatial density, enabling the visualization of sensor data in 3D space. Additionally, to support low-latency visualization of dynamic sensor data, we design latent reuse caches at each step, enabling just-in-time AR presentations (Figure 9).
@@ 6.2. AR Production
From the CLIP embeddings and sensor context, we produce 2D visualizations. Building on this foundation, we extend these images into 3D AR content. Traditional 3D content generation typically uses mesh-based models that represent only surface structures, leaving interiors empty. This approach is not suitable for sensor data, which often encapsulates volumetric phenomena. For example, environmental readings like temperature or humidity demand a representation that is volumetrically consistent rather than a hollow shell.
To address this limitation, we develop a 3D generation approach that incorporates sensor embeddings and manifestations to produce volumetric models that accurately reflect the spatial context of sensor data. Leveraging 3DGS (Kerbl et al., 2023; Liu et al., 2023; Tang et al., 2023), we generate spatially-aware visualizations that capture the volumetric and spatial aspects of sensor data, such as temperature distributions within a space or air quality surrounding a street.
Our method integrates sensor embeddings and manifestations into the 3D generation pipeline by adapting the DreamGaussian (Tang et al., 2023) framework. We use sensor-specific embeddings and 2D images as inputs to generate volumetric representations that preserve the spatial characteristics of the sensor data. To create truly volumetric visualizations, we introduce a density-aware generation technique that applies controlled random dropout to Gaussian points, creating transparent areas that allow the model to fill both internal and external spaces. The dropout rate is determined by the sensor schema and readings through a density encoder, ensuring alignment with the sensor data’s characteristics. This design ensures the generated AR content’s interior is fully populated with Gaussian points.
@@
Embedding heterogeneous sensor data into visual spaces poses significant challenges, particularly in extreme conditions. Our proposed FM-assisted pipeline addresses these issues by leveraging large language models (LLMs) and schema generation. LLMs offer a robust solution for addressing the complexity and diversity of sensor data. These models excel in processing and understanding the intricate relationships between various data modalities. A key feature of our pipeline is the creation of a sensor schema that includes anchors specific to each sensor type and context. This schema is constructed using a combination of data collection, domain knowledge, and the capabilities of LLMs, enabling zero-shot interpretation by generating meaningful visual manifestations without prior task-specific training.
Our pipeline is specifically designed to handle sensor data using zero-shot learning capabilities, which allows for the automatic processing of new sensor inputs. We construct a schema for each sensor type that anchors the sensor data to its contextual meaning, leveraging LLMs for schema refinement. With careful prompt design, LLMs automatically generate sensor metadata descriptions, which are then visualized. Additionally, advanced image generation models are employed to create detailed and accurate visual constraints, ensuring consistency with the defined sensor schema and manifestation.
@@ 6.3. Latent Reuse for Fast Generation
The generation of both images and 3D Gaussian Splatting is computationally expensive. Even with high-end GPUs like the Nvidia RTX 4090, creating a single image can take over 5 seconds per frame, while generating a 3D scene with DreamGaussian (Tang et al., 2023) may require up to 50 seconds. These time demands significantly impact the practicality of sensor visualizations, especially when dealing with diverse and rapidly changing data. To mitigate this, we designed a latent reuse caching mechanism that accelerates the generation of both images and 3D scenes, achieving a 7×7$\times$7 × reduction in processing time without compromising output quality.
@@ 7.3. Evaluations of Visualization Quality
Figure 11. Visualization of human emotional states.
Figure 12. Environmental data visualization – temperature and humidity.
Figure 13. Model performance comparison across contexts (left) and user survey considerations (right).
@@
To assess the quality of generated content, we conducted a survey with 565 participants recruited through social media and forums, a total of 485 valid responses, of which 91.2% had prior experience with sensor technologies, enhancing their evaluation accuracy. The survey aimed to compare Vivar with state-of-the-art image generation models based on three criteria, each rated on a scale from 1 to 5, where 1 is bad, 3 is okay, and 5 is awesome:
@@
    Environment: Participants imagined being in a room with temperature and humidity sensors. The visualizations were intended to reflect these changes.
@@
    Emotion: Participants viewed images depicting a person experiencing a range of emotions (Figure 13).
@@
    Entertainment: Participants imagined listening to different types of music, from soft to noisy, expecting the visualizations to correspond to these auditory variations.
@@
ATitan 0.65 0.68 0.64 0.66
DallE3 0.62 0.70 0.78 0.70
SDE 0.53 0.60 0.60 0.58
SDP 0.59 0.48 0.53 0.53
Vivar (Ours) 0.83 0.72 0.83 0.79
Improvement 0.18 (27.7%) 0.02 (2.9%) 0.05 (6.4%) 0.09 (12.9%)
Table 2. Performance comparison across different evaluation metrics. The Improvement row shows the absolute and relative (%) difference between Vivar and the second-best model for each metric.
@@
Results and Analysis
Table 2 summarizes the processed scores for each model, with the highest scores for each metric highlighted in green and the lowest in red. To handle variation in rating preference, we applied response style adjustment through baseline normalization, subtracting the minimum score within each question-factor group per participant. This preprocessing preserved relative differences between ratings while establishing a common reference point. Our proposed system, Vivar, achieved the highest scores across all metrics. The most notable improvement was in coherence (0.83, 0.18 higher than the second-best model), attributed to the continuity provided by embedding blending, which results in more logically consistent image generation. Vivar also excelled in faithfulness (0.72), reflecting the effectiveness of our manifestation generation in accurately representing sensor data. For sensitivity, Vivar (0.83) outperformed DALL-E 3 (0.78), despite the latter’s strong language understanding capabilities, validating our design goal that visualizations accurately reflect changes in values. Overall, Vivar achieved the highest performance score (0.79), outperforming DALL-E 3 and ATitan by approximately 19.7% and 12.9% respectively, when accounting for user-prioritized criteria.
Multi-modality Handling
Figure 13 compares our barycentric interpolation approach with the baseline (using text prompts with Stable Diffusion) and inverse distance weight (IDW) interpolation. Our method correctly produces visualizations with appropriate semantics while maintaining consistency across the sensor space. When transitioning from cold & dry to cold & humid conditions, IDW interpolation fails to preserve the semantic information of cold weather, as it simply averages the embeddings and cannot preserve semantic correctness (circled in red). A user study involving an additional 18 participants who watched 386 frames of continuous visual productions showed unanimous agreement (100%) that our approach provided superior coherence and faithfulness in visualization. All participants indicated a clear preference for our method over the baseline and IDW interpolation.
@@ 9   Conclusion
In this work, we addressed the challenge of visualizing multi-modal and context-dependent sensor data by introducing a novel cross-modal embedding method that blends sensor readings into a pre-trained visual embedding space, guided by the shading principle. By incorporating diverse manifestations and 3DGS, our approach enables accurate, consistent visual representations with smooth transitions between data points, maintaining semantic coherence. The inclusion of a latent reuse cache further enhances adaptability to dynamic sensor readings. Through evaluations involving 503 participants, we demonstrated that our solution outperforms state-of-the-art visual generation models in consistency and representativeness, making it better suited to sensor data visualization. A user study with domain experts further validated the system’s accuracy, reliability, and practical applicability across diverse downstream tasks. This work establishes a foundation for intuitive, coherent, and unbiased sensor data visualization by bridging domains. Drawing inspiration from graphics and shading principles, we combine large pre-trained models with explainable transformations, opening new avenues for advancing cross-modal embedding techniques.